# Autodeleveraging: Impossibilities and Optimization

### v1

@@ 8.3 Incentive Compatibility and Constraints
We frame the problem as a Stackelberg game: the exchange (leader) commits to a policy, and traders (followers) best-respond. Follower Incentive Compatibility requires traders’ actions—liquidity provision and liquidations—to satisfy feasibility and non-bankruptcy conditions given this commitment. The exchange’s policy is subject to Revenue Constraints (ensuring profitability via fees and net liquidation revenues) and Solvency Constraints (bounding tail risk via coherent measures like ES or CVaR [RockafellarUryasev2002, BCBS2019FRTB]). These constraints define the feasible region for the loss minimization problem.
8.4 ADL Constraints and Objective
Composite Loss Function.
We formulate the ADL control problem as minimizing a composite loss function that explicitly balances exchange revenue, solvency, and fairness. Let $\pi_{t}=(\theta_{t},h_{t})$ denote the ADL policy at round $t$, where $\theta_{t}\in[0,1]$ is the severity (fraction of deficit socialized) and $h_{t}\in[0,1]^{n}$ parameterizes the inner allocation rule (e.g., weights for RAP or Pro-Rata). Based on our empirical framework, the per-round loss $\ell_{t}(\pi_{t})$ is defined as:
$\ell_{t}(\pi_{t})\;=\;\lambda_{\text{rev}}\cdot\mathcal{L}^{\text{rev}}_{t}(\pi_{t})\;+\;\lambda_{\text{solv}}\cdot\mathcal{L}^{\text{solv}}_{t}(\pi_{t})\;+\;\lambda_{\text{fair}}\cdot\mathcal{L}^{\text{fair}}_{t}(\pi_{t}),$
where the components are:
  • Revenue Loss ($\mathcal{L}^{\text{rev}}$): The discounted stream of future fees lost due to liquidating positions. This incentivizes the exchange to preserve valuable, high-volume users when possible.
  • Solvency Cost ($\mathcal{L}^{\text{solv}}$): A penalty on the residual deficit (the portion of the loss $D_{t}$ not covered by the ADL recovery) and overshoot (excessive liquidation beyond what is needed). This ensures the mechanism actually restores system health without unnecessary destruction.
  • Fairness Cost ($\mathcal{L}^{\text{fair}}$): Penalties based on the Profit-to-Solvency Ratio (PTSR) and Profit-Margin Ratio (PMR). These terms enforce that the burden of ADL is distributed equitably relative to traders’ profitability and risk contribution, preventing the “death spiral” of adverse selection.
@@ 8.4 ADL Constraints and Objective
We model the interaction between the exchange and traders as a Stackelberg game where the exchange (leader) commits to a policy $\pi_{t}$, and traders (followers) subsequently respond (e.g., by adjusting liquidity or closing positions). This commitment power is crucial. As we show in Appendix F.3 (Proposition F.6), a simultaneous-move (Nash) game admits a “bad” equilibrium where traders withdraw liquidity anticipating high severity, which in turn forces the exchange to set high severity to cover the deficit. In the Stackelberg model, the exchange can commit to a policy that eliminates this inefficient equilibrium, coordinating the market towards a high-liquidity, lower-severity outcome that improves welfare for all participants.
@@ 8.5 Learning the Optimal ADL Policy
    Choose a mirror map $\psi$ on the policy domain $\Pi$ (e.g., a separable entropy or a Bregman geometry aligned with capped pro–rata).
@@
Why MDIC vs. Online Convex Optimization.
All constraints in $\mathcal{C}_{t}^{\mathrm{IC}}$ are convex but can be non-smooth (e.g. hinge losses, maximizing over scenarios). Mirror descent handles this composite structure while matching the geometry: KL geometry on the simplex reduces diameter to $O(\log|W_{t}|)$; the log-barrier keeps $\theta$ interior and gives stable 1-D projections. This improves constants relative to Euclidean OGD while preserving the same rate.
Algorithmic Guarantees.
Assume $f_{t}$ is $\ell$-Lipschitz in the dual norms of $(\phi,\Phi)$, the follower is an $\varepsilon$-best responder, and $\mathcal{C}_{t}^{\mathrm{IC}}$ is nonempty with bounded diameter in the same norms. Then MDIC with $\eta_{t}\propto 1/\sqrt{T}$ achieves
@@
and enforces the IC rows per round (or, in a primal–dual variant with long-term ES, sublinear average violation). A full theorem and proof (via the standard MD telescoping bound plus feasibility of the Bregman projection) appear in Appendix F.2; see also Theorem F.2 for the unconstrained MD bound. Under these convexity, Lipschitz, and feasibility assumptions, MDIC attains $O(\sqrt{T})$ regret and sublinear average IC/solvency violations, whereas any static policy incurs $\Omega(T)$ regret on an alternating environment (see Appendix F.2 for details).
Linear Regret for Static Strategies.
Consider a simplified environment that alternates between two regimes: $A$ (liquidity) and $B$ (stress). We fix a maximum severity cap $\bar{\theta}$ (e.g., $\bar{\theta}=0.30$). The per-round loss function $f_{t}(\theta)$ captures the trade-off between revenue and solvency:
[…]
The optimal static policy must choose a fixed $\theta\in[0,\bar{\theta}]$. However, any fixed $\theta$ incurs linear regret:
$\sum_{t=1}^{T}(f_{t}(\theta)-f_{t}(\theta^{\star}_{t}))=\frac{T}{2}\theta+\frac{T}{2}(\bar{\theta}-\theta)=\frac{T}{2}\bar{\theta}=\Omega(T).$
In contrast, MDIC adapts $\theta_{t}$ online. Using a standard mirror descent bound (Theorem F.2), MDIC achieves
$\mathsf{Regret}_{T}=O(\bar{\theta}\sqrt{T}).$
Numerically, for $T=10^{4}$ and $\bar{\theta}=0.30$, a static policy suffers regret $\approx 1500$, while MDIC suffers $\approx 30$.
@@ 8.6 Follower Robustness
So far, we have focused on the leader (e.g. the exchange), as they have to commit to a policy. However, it is possible that if the follower (i.e., a trader who reduces the deficit) knows enough about the leader’s strategy to adversely select against other traders. We address the strategic behavior of traders (“followers”) in response to ADL policies, focusing on adverse selection and timing games that naturally occur. Our coverage of this will be minimal with most details in Appendix F.4.
@@
If an ADL policy socializes losses uniformly (like Pro-Rata), low-risk traders subsidize high-risk traders. If this subsidy exceeds the utility of trading, low-risk traders exit, increasing both the exchange’s average risk and probability of future deficits. This is a direct consequence of the follower’s strategic response: low-risk traders, observing a policy that penalizes them disproportionately, rationally choose to exit (or not participate), leaving the exchange with a riskier pool of traders.
Formally, a participation constraint $\mu_{i}-\mathop{\bf E{}}[H_{\pi}(\lambda_{i},\mathcal{P}_{n})]\geq u_{0}$ must hold. Here $\mu_{i}$ denotes trader $i$’s baseline per-round utility in the absence of ADL haircuts, $H_{\pi}(\lambda_{i},\mathcal{P}_{n})$ is the random haircut mass assigned to a type-$i$ winner with leverage $\lambda_{i}$ under policy $\pi$ when the market state is $\mathcal{P}_{n}$, and $u_{0}$ is the reservation utility defined in Appendix F.4.1. Under Pro-Rata, $\mathop{\bf E{}}[H_{\pi_{PR}}]$ depends on average leverage, which penalizes low-risk traders and potentially causes a “death spiral” of exits. This “death spiral” refers to the adverse selection that the exchange faces as the pool of traders becomes increasingly risky from a leverage standpoint.
In contrast, RAP’s haircut is proportional to risk weight $\lambda_{i}g(\lambda_{i})$, minimizing the burden on low-risk traders and stabilizing the pool. Appendix F.4.1 formalizes this participation test and provides explicit death-spiral examples.
@@
Traders may act as “Backstop Liquidity Providers” (BLP) by voluntarily filling liquidation orders to reduce the deficit $D_{t}$. For instance, traders who deposit to Hyperliquid’s HLP pool around the time of an ADL event do this. This presents an optimal stopping problem: a BLP can absorb the deficit immediately (incurring market risk and execution costs) or wait, hoping that the exchange reduces ADL severity $\theta_{t}$ (and thus the potential haircut) in subsequent rounds. If the exchange employs a policy like exponential backoff that lowers severity over time, it incentivizes traders to wait, potentially exacerbating the deficit.
To prevent such strategic delays, the exchange must satisfy a “No-Wait” constraint: $\theta_{t}D_{t}\geq\Gamma_{t}$, where $\Gamma_{t}$ is the liquidity premium. Intuitively, $\Gamma_{t}$ quantifies the cost of immediate intervention, capturing the execution slippage and inventory risk premium required for a BLP to step in. By adjusting incentives that are paid to BLPs, an exchange can incentivize “no-wait” conditions and lower the overall time to recover solvency. We formalize this with the following proposition:
@@
Let $\Gamma_{t}$ be the liquidity premium required for a Backstop Liquidity Provider (BLP) to absorb deficit $D_{t}$ immediately rather than waiting for $\Delta t$. If the severity policy $\theta_{t}$ is a decreasing function of time (e.g., exponential backoff), a BLP will withhold liquidity unless the immediate haircut cost is cheaper than the discounted future cost:
@@
Appendix F.4.2 details the construction of $\Gamma_{t}$ and proves this proposition.
@@ 8.7 Dynamic Phase Transition in ADL
In Section 5, we established that static ADL rules face fundamental scaling limits. To quantify the efficiency gap between static (Nash) and dynamic (Stackelberg) control, we evaluate the system using welfare functionals $W(\pi)$ corresponding to the trilemma (Proposition I.7) objectives. We consider both a fairness welfare $W_{\text{Fair}}(\pi)=\mathsf{PTSR}_{T}(\pi)$ and a revenue welfare $W_{\text{LTV}}(\pi)=\text{LTV}_{T}(\pi)$. The severity load $\kappa_{n}=\theta_{n}n/b_{n}$ serves as the order parameter for the fairness transition, while the heavy-tail index $\alpha$ (see §3.1.2) governs the revenue transition. Together, they drive a phase transition in the efficiency of static policies relative to dynamic ones.
Why PoA for Nash vs. Stackelberg?
PoA is the standard metric for quantifying efficiency loss from strategic behavior across equilibrium concepts [Roughgarden2005SelfishRouting, roughgarden15intrinsic]. In Stackelberg settings, a leader’s commitment reshapes follower best responses and can systematically lower inefficiency relative to simultaneous-move (Nash) play. This is well documented in dynamic game models [BasarOlsder1999] and in smoothness-based PoA analyses that extend uniformly across equilibrium notions, including sequential and composed mechanisms [SyrgkanisTardos2013, roughgarden15intrinsic]. In our ADL context, the venue is a natural leader that can commit to a severity path, while traders react myopically to current and anticipated haircuts. This makes a PoA comparison between one-shot (Nash-like) and multi-round (Stackelberg) control both natural and informative.
We define the Price of Anarchy (PoA) as the ratio of the optimal dynamic Stackelberg welfare to the welfare of a static, one-shot policy $\pi^{\mathrm{stat}}_{n}$:
@@
Using smoothness arguments [roughgarden15intrinsic] and extreme-value analysis, we obtain PoA bounds that are robust to the precise equilibrium concept. Appendix G formalizes our mean-field scaling assumptions and presents the theorem. Informally:
@@
  • Bounded PoA (Static Sufficient): If severity is EV-scaled ($\sup\kappa_{n}<\infty$), leverage is balanced ($\ell_{n}^{+}\asymp\ell_{n}^{-}$), and impact is smooth ($\mu_{\mathrm{OB}}<1$), then static ADL is constant-factor optimal for both Fairness and Revenue: $\sup_{n}\mathrm{PoA}^{\text{Fair}}_{n}<\infty$ and $\sup_{n}\mathrm{PoA}^{\text{LTV}}_{n}<\infty$.
  • Unbounded PoA (Dynamic Necessary): If severity is excessive ($\kappa_{n}\to\infty$) or leverage is concentrated ($\ell_{n}^{+}/\ell_{n}^{-}\to 0$), then static policies fail on at least one objective: $\max(\mathrm{PoA}^{\text{Fair}}_{n},\mathrm{PoA}^{\text{LTV}}_{n})\to\infty$. Specifically, heavy tails force static policies to choose between fairness collapse ($\theta\approx 1$) or revenue collapse ($\delta\approx 1$).
@@
Practically, this result suggests that exchanges can screen assets using historical data to determine whether a simple static rule suffices or if a dynamic policy is required. Formal statements and proofs are given by Theorem G.3 and Proposition G.5 in Appendix G, and by Example G.4.
9 Empirical Analysis: The October 10 Event
We evaluate the proposed ADL mechanisms using a high-fidelity replay of the October 10, 2025 liquidation cascade on Hyperliquid [CoinDesk2025LargestLiquidations]. This event serves as a practical testbed for the theoretical phase transitions identified in Section 8.7, characterizing the performance gap between static and dynamic policies in a high-stress regime ($D_{t}\gg 0$).

### v3

@@ 8.3 Incentive Compatibility and Constraints
We frame the problem as a Stackelberg game: the exchange (leader) commits to a policy, and traders (followers) best-respond. We impose follower incentive compatibility (FIC) and leader-side revenue and solvency constraints (LIC) to ensure the policy is feasible for traders and sustainable for the venue.
Follower Incentive Compatibility (FIC).
Winners must remain solvent and within haircut caps:
$0\;\leq\;h_{t,i}\;\leq\;\beta_{t,i},\qquad(1-h_{t,i})e_{t,i}^{+}\;\geq\;\underline{e}_{t,i},$
where $\beta_{t,i}$ enforces per-account haircut limits and $\underline{e}_{t,i}$ enforces minimum post-ADL equity (cf. (26)–(27)). These bounds guarantee feasibility (no negative winners) and rule out haircut paths that would violate venue-level haircut caps (as we saw in §2.4).
Leader Incentive Compatibility (LIC).
Revenue must stay above a floor and solvency above a safety threshold:
$\mathop{\bf E{}}[\mathrm{LTV}_{t}]\;\geq\;(1-\epsilon)\,\mathrm{LTV}_{\text{pre}},\qquad\mathsf{IF}_{t}\;\geq\;\delta\;\;\text{or}\;\;\mathrm{CVaR}_{\alpha}(-\mathsf{IF}_{t})\;\leq\;\kappa.$
These ensure the policy is not revenue-destructive (per the churn model in §9) and keeps the exchange solvent both on-path and in the tails (§2.3). These constraints define the feasible set for any policy path $\{\theta_{t},h_{t}\}$; we defer the full optimization program to Appendix G.3.
@@
We model the interaction between the exchange and traders as a Stackelberg game where the exchange (leader) commits to a policy $\pi_{t}$, and traders (followers) subsequently respond (e.g. by adjusting liquidity or closing positions). This commitment power is crucial. As we show in Appendix G.3 (Proposition G.6), a simultaneous-move (Nash) game admits a “bad” equilibrium where traders withdraw liquidity anticipating high severity, which in turn forces the exchange to set high severity to cover the deficit, often violating the LIC floor on revenue. In the Stackelberg model, the exchange commits to a policy that satisfies FIC/LIC and rules out this inefficient equilibrium, coordinating the market toward a high-liquidity, lower-severity outcome that keeps both trader feasibility and venue sustainability intact.
@@ 8.4 Learning the Optimal ADL Policy
Given these constraints, we can treat severity and haircuts as an online control problem: choose $\{\theta_{t},h_{t}\}$ after each price shock, conditioned on past shocks and responses, to minimize loss while respecting FIC/LIC. Because losses admit convex surrogates, we can apply classical online schemes (e.g. mirror descent) and enforce constraints via primal–dual updates so FIC/LIC remain satisfied at each round. We use mirror descent because it naturally handles convex losses with simple geometry (RAP/Pro-Rata) and pairs cleanly with primal–dual updates, matching the Stackelberg setting of committing to a policy while respecting per-round constraints [62, 63, 118, 103, 66].
@@
    Choose a mirror map $\psi$ on the policy domain $\Pi$ (e.g. a separable entropy or a Bregman geometry aligned with capped pro–rata).
@@
Why MDIC.
The IC set is convex but non-smooth (e.g. caps, max over scenarios). Mirror descent matches this geometry (KL on the simplex, log-barrier on $\theta$) and keeps updates stable while preserving $O(\sqrt{T})$ rates.
Algorithmic Guarantees and Why Regret Matters.
If $f_{t}$ is Lipschitz, the follower is $\varepsilon$-best-response, and the IC set has bounded diameter, MDIC with $\eta_{t}\propto 1/\sqrt{T}$ yields
@@
with per-round IC feasibility (or sublinear average violation in the long-term CVaR variant). We care about regret because it benchmarks any adaptive policy against the best dynamic policy in hindsight. Showing that MDIC achieves $O(\sqrt{T})$ regret means MDIC learns fast, while $\Omega(T)$ regret shows static rules (Queue, RAP, Pro-Rata) can be worst-case bad. Proofs are in Appendix G.2.
Static Policies Have Linear Regret.
In an alternating two-regime example (liquidity vs. stress) any fixed $\theta$ incurs $\Omega(T)$ regret, whereas MDIC adapts and achieves $O(\sqrt{T})$ (proven in Appendix G.2). This captures why static Queue, RAP, Pro-Rata are fragile in adversarial environments.
@@ 8.5 Follower Robustness
So far, we have focused on the leader (e.g. the exchange), as they have to commit to a policy. However, it is possible that if the follower (i.e., a trader who reduces the deficit) knows enough about the leader’s strategy to adversely select against other traders. We address the strategic behavior of traders (“followers”) in response to ADL policies, focusing on adverse selection and timing games. Our description of follower behavior will be minimal with most details in Appendix G.4.
@@
As we saw earlier, if an ADL policy socializes losses uniformly without regards to risk (like Pro-Rata), low-risk traders subsidize high-risk traders. If the cost of subsidy exceeds the utility of trading for low-risk traders, they exit. This increases both the exchange’s average risk and probability of future deficits and can be viewed as a form of adverse selection. This is a direct consequence of the follower’s strategic response: low-risk traders, observing a policy that penalizes them disproportionately, rationally choose to exit (or not participate), leaving the exchange with a riskier pool of traders.
Let $\mu_{i}$ be the baseline utility of trader $i$ and let $H_{\pi}(\lambda_{i}\mathcal{P}_{n})$ be the random haircut variable for a trader with leverage $\lambda_{i}$. The "utility is greater than subsidy" condition can be described as $\mu_{i}-\mathop{\bf E{}}[H_{\pi}(\lambda_{i},\mathcal{P}_{n})]\geq u_{0}$, where $u_{0}$ is a baseline minimum utility a low-risk trader accepts. Under Pro-Rata, $\mathop{\bf E{}}[H_{\pi_{PR}}]$ depends on the average leverage $\overline{\lambda}=\frac{1}{n}\sum_{i}\lambda_{i}$, which penalizes low-risk traders with $\lambda_{i}<\overline{\lambda}$. The worst case is if there is a ‘death spiral’ of exits, where low-risk traders leaving sequentially increase the probabilty of other traders leaving, as the average $\overline{\lambda}$ increases on each exit. This is adverse selection that the exchange faces as the pool of traders becomes increasingly risky from a leverage standpoint. We formalize this in Appendix G.4.1 and provide explicit examples of a death spiral.
@@
Traders may act as “Backstop Liquidity Providers” (BLP) by voluntarily filling liquidation orders to reduce the deficit $D_{t}$. For instance, traders who deposit to Hyperliquid’s HLP pool around the time of an ADL event act as BLPs. This presents an optimal stopping problem: a BLP can absorb the deficit immediately (incurring market risk and execution costs) or wait, hoping that the exchange reduces ADL severity $\theta_{t}$ (and thus the potential haircut) in subsequent rounds. If the exchange employs a policy like an exponential backoff that lowers severity over time, it incentivizes traders to wait, potentially exacerbating the deficit. To prevent such strategic delays, the exchange needs to provide a financial incentive to ensure that BLPs add liquidity in a timely manner.
This incentive often comes in the form of yield premium $\Gamma_{t}$. This premium can be viewed as an incentive (e.g., a rebate or a yield) that the exchange must pay to compensate BLPs for the risk of absorbing the deficit early. Intuitively, $\Gamma_{t}$ quantifies the cost of immediate intervention, capturing the execution slippage and inventory risk premium required for a BLP to step in. By adjusting these incentives, an exchange can encourage "no-wait" conditions to speed up solvency recovery. Intuitively, these “No-Wait” conditions correspond to solvency-like constraints: $\theta_{t}D_{t}\geq\Gamma_{t}$. We formalize this with the following proposition:
@@
Let $\Gamma_{t}$ be the liquidity premium required for a Backstop Liquidity Provider (BLP) to absorb deficit $D_{t}$ immediately rather than waiting for $\Delta t$. If the severity policy $\theta_{t}$ is a decreasing function of time (e.g. exponential backoff), a BLP will withhold liquidity unless the immediate haircut cost is cheaper than the discounted future cost:
@@
Appendix G.4.2 details the construction of $\Gamma_{t}$ and proves this proposition.
@@ 8.6 Dynamic Phase Transition in ADL
In Section 5, we established that static ADL rules face fundamental scaling limits. However, our results in that section were qualitative in nature, making it hard to use those results to design better mechanisms. One question that one might have is if there is a way to classify, based on properties of an asset (e.g. volatility, liquidity, etc.) whether one can use a simple static ADL strategy or if one needs a full dynamic strategy. Here we will construct conditions for quantifying an answer to this question.
To quantify the efficiency gap between static (Nash) and dynamic (Stackelberg) control, we evaluate the system using welfare functions $W(\pi)$ corresponding to the trilemma (Proposition J.7) objectives. Welfare functionals quantify how fair, profitable, or solvent the entire system is (as opposed to individual traders). Specifically, we show there is a phase transition that depends on the liquidity and volatility of the asset pair that determines when one needs a dynamic methods (like MDIC) or if one can use a simpler algorithm (like RAP). The phase transition is defined by an order parameter, which is a dimensionless quantity that captures the relative strength of the ADL severity to the market’s liquidity and volatility. This phase transition occurs in the efficiency gap between static and dynamic methods — i.e., how much better at preserving solvency and/or revenue that a dynamic method is over static method.
Price of Anarchy for Nash vs. Stackelberg.
The Price of Anarchy (PoA) is the standard metric for quantifying efficiency loss from strategic behavior across equilibrium concepts [100, 101]. In Stackelberg settings, a leader’s commitment reshapes follower best responses and can systematically lower inefficiency relative to simultaneous-move (Nash) play. This is well documented in dynamic game models [18] and in smoothness-based PoA analyses that extend uniformly across equilibrium notions, including sequential and composed mechanisms [112, 101]. In our ADL context, the venue is a natural leader that can commit to a severity path, while traders react myopically to current and anticipated haircuts. This makes a PoA comparison between one-shot (Nash-like) and multi-round (Stackelberg) control both natural and informative.
We define the PoA for a welfare function $W$ as the ratio of the optimal dynamic Stackelberg welfare to the welfare of a static, one-shot policy $\pi^{\mathrm{stat}}_{n}$:
@@
We consider both a fairness welfare $W_{\text{Fair}}(\pi)=\mathsf{PTSR}_{T}(\pi)$ and a revenue welfare $W_{\text{LTV}}(\pi)=\text{LTV}_{T}(\pi)$. The severity load $\kappa_{n}=\theta_{n}n/b_{n}$ serves as the order parameter for the phase transition in fairness, while the heavy-tail index $\alpha$ (see §3.1.2) governs the phase transition in revenue. Using extreme-value scaling and mean-field equilibrium assumptions, we obtain PoA bounds that are robust to the precise equilibrium concept. Appendix H formalizes the scaling assumptions and presents the theorem. Informally:
@@
  • Bounded PoA (Static Sufficient): If the severity load stays bounded ($\sup_{n}\kappa_{n}<\infty$), leverage masses are comparable ($\ell_{n}^{+}\asymp\ell_{n}^{-}$), and deficits are light-tailed, then static ADL is constant-factor optimal: $\sup_{n}\mathrm{PoA}^{\text{Fair}}_{n}<\infty$ and $\sup_{n}\mathrm{PoA}^{\text{LTV}}_{n}<\infty$.
  • Unbounded PoA (Dynamic Necessary): If the load diverges ($\kappa_{n}\to\infty$) or leverage is concentrated ($\ell_{n}^{+}/\ell_{n}^{-}\to 0$), Fairness PoA blows up. If deficits are heavy-tailed ($\alpha<2$) and solvency is enforced, Revenue PoA also diverges. Static policies must sacrifice at least one objective, while a dynamic controller can maintain both.
@@
Practically, this result can be used to provide simple heuristics for screening assets using historical data to determine whether a simple static rule suffices or if a dynamic policy is required. Formal statements and proofs are given by Theorem H.3 and Proposition H.5 in Appendix H, and by Example H.4.
9 Empirical Analysis under Partial Observability: The October 10 Event
Partial Observation model.
Our reconstruction uses publicly available fills, marks, and account-level aggregates. It is not a full execution and clearing ledger: complete settlement logic, internal netting, fee and buffer accounting, and some timing details are unobserved. These observables are sufficient to compute our empirical targets (deficits and winner-side PNL capacity) under explicit definitions. Accordingly, strict accounting equalities are treated as diagnostics rather than identities. Throughout this section, “counterfactual” refers to counterfactuals under this public-data observation model, not to a complete ledger identity.
We evaluate the proposed ADL mechanisms using a reconstruction [^13] We thank Mauricio Trujillo (@conejocapital) and Hydromancer for providing, cleaning, and validating this dataset. of the October 10, 2025 liquidation cascade on Hyperliquid [99]. We use this replay to reconstruct production wealth removed from winners via a two-pass no-ADL counterfactual, compute the instantaneous bankruptcy-gap proxy $B_{t}^{\mathrm{needed}}$, and compare production to transparent benchmark allocations under a profits-only (PNL) haircut-capacity constraint.